A discount store sold plastic cups for $2.75 each and ceramic cups for $5.50 each.
If 500 cups were sold for a total of $1897.50, how many cups of each type were sold?

Answer by mananth(16946)

plastic cups x numbers
ceramic cups y numbers

1 x + 1 y = 500 .............1
2.75 x + 5.50 y = 1897.50 .............2
Eliminate y
multiply (1)by -5.5
Multiply (2) by 1
-5.5 x -5.5 y = -2750
2.75 x + 5.5 y = 1897.5
Add the two equations
-2.75 x = -852.50
/ -2.75
x = 310
plug value of x in (1)
1 x + 1 y = 500
310 + 1 y = 500
1 y = 500
1 y = 190
y = 190
plastic cups numbers 310
ceramic cups numbers 190
m.ananth@hotmail.ca